# You Only Look at Once for Real-time and Generic Multi-Task

### v1

@@ IV-B2 Multi-task Result
This part presents the results on multi-task experiment, encompassing object detection, drivable area segmentation, and lane line segmentation.
We follow the YOLOP settings in the detection task, combining the car, bus, truck, and train into a “vehicle” classification. The comparison results are shown in Table II. Based on the quantitative results, we can see that both versions of our model achieve the best performance in terms of mAP50. This indicates that our model’s accuracy is excellent in predicting detected targets. Especially when compared to MultiNet and Faster R-CNN. Additionally, A-YOLOM(n) has a complexity level for its backbone, detection neck, and head that is comparable to YOLOv8n(det). However, A-YOLOM(n) exhibits significant improvement over YOLOv8n(det) in both Recall and mAP50. This demonstrates that in multi-task learning, various tasks can implicitly assist and further enhance the performance of individual tasks. On another note, YOLOv8n(det) outperforms YOLOv5s because the ’s’ scale model usually has a more complex backbone to enhance performance. One issue with our model is that the Recall is not satisfactory. YOLOP and DLT-Net achieve better Recall than our model. We believe this is due to the high weight of $\lambda_{box}$ in $\mathcal{L}_{det}$. This means the model is more conservative and focuses on achieving a higher mAP at the expense of Recall performance. We believe that mAP50 better reflects the comprehensive performance in the detection task. Besides, our model achieves real-time performance, whereas YOLOP cannot when the batch size is set to 1 according to Table I.
@@
Model
Recall (%)
mAP50 (%)
MultiNet
81.3
60.2
DLT-Net
89.4
68.4
Faster R-CNN
81.2
64.9
YOLOv5s
86.8
77.2
YOLOv8n(det)
82.2
75.1
YOLOP
88.6
76.5
A-YOLOM(n)
85.3
78.0
A-YOLOM(s)
86.9
81.1
For the drivable area segmentation task, Table III provides the quantitative results. Our model achieved the second and third best performances in terms of mIoU. YOLOP outperforms our model because they customized the loss function for the segmentation task. In terms of mIoU, A-YOLOM(n) and A-YOLOM(s) trail YOLOP by 1% and 0.5% respectively. We believe this result is acceptable. While our model might exhibit a slight loss in performance, it is more flexible and faster. Additionally, our model performs much better than other models, such as YOLOv8n(seg), MultiNet, and PSPNet. It’s worth noting that while YOLOv8n(seg) is faster than our model, its deployment cost is higher, and its performance is significantly inferior to ours.
@@
Model
mIoU (%)
MultiNet
71.6
DLT-Net
72.1
PSPNet
89.6
YOLOv8n(seg)
78.1
YOLOP
91.6
A-YOLOM(n)
90.5
A-YOLOM(s)
91.0
For the lane line segmentation task, Table IV provides the quantitative results. A-YOLOM(s) achieves the best performance in terms of both accuracy and IoU. Specifically, it delivers competitive results in accuracy and is 2.6% higher in IoU compared to YOLOP. It’s worth noting that our model maintains the same structure and loss function across all segmentation tasks, eliminating the need for adjustments when encountering a new segmentation task. Additionally, YOLOv8 (seg) is slightly inferior to ours. Based on results in Table III and Table IV, we believe that our proposed neck and head are better suited for the segmentation task compared to YOLOv8 (seg). Since we use balanced accuracy for evaluation, Enet, SCNN, and ENet-SAD employ different accuracy calculation approaches. Therefore, we cannot directly compare our accuracy with theirs. However, our results in IoU metrics are significantly better than theirs. This alone sufficiently demonstrates the superior performance of our model compared to theirs.
@@
Model
Accuracy (%)
IoU (%)
Enet
N/A
14.64
SCNN
N/A
15.84
ENet-SAD
N/A
16.02
YOLOv8n(seg)
80.5
22.9
YOLOP
84.8
26.5
A-YOLOM(n)
81.3
28.2
A-YOLOM(s)
84.9 28.8
@@ IV-B3 Visualization
This section presents a visual comparison between YOLOP and our model. We evaluate performance not only in favourable weather conditions but also under adverse conditions, including strong sunlight, nighttime, rain, and snow. We will analyze each scene in turn.
Figure 2 displays the results from a sunny day. As we know, strong sunlight can affect a driver’s vision. Similarly, it impacts the model’s performance. This poses challenges for the drivable area and lane line segmentation performance. In this challenging scenario, our model outshines YOLOP. Specifically, under the conditions of intense sunlight (as seen in the left and right images), our model delivers accurate lane line predictions and provides smoother indications of drivable areas. Additionally, our model exhibits greater accuracy in detecting smaller and farther vehicles. As observed in the middle and right images, our model successfully detects vehicles located far in the distance, both on the road in opposite driving directions and near the house. We believe that our model outperformed YOLOP in this scene according to the visualization results.
@@
Figure 3 displays the results from night scenes. In nighttime scenes, the image quality decreases due to limited lighting and glare from oncoming vehicles. This poses a challenge for the model to make accurate predictions. Under this challenge, our model consistently produces more accurate and smoother predictions for both lane lines and drivable areas. A-YOLOM(s) excels not only in the segmentation task but also in the detection task. Specifically, it surpasses A-YOLOM(n) by accurately detecting vehicles travelling in the opposite direction, even amidst glare, and holds its own against YOLOP. Based on the detection results from the right image, YOLOP slightly outperforms our model in detecting distant vehicles at night. Nonetheless, our model produces significantly better segmentation results compared to YOLOP. Especially in the middle image, YOLOP mistakenly predicts the opposite lane as a drivable area. Such mispredictions are extremely hazardous for autonomous driving tasks.
@@
Figure 4 displays the results from a rainy day. Rain will increase road surface reflection, thereby impacting a driver’s judgment. This issue also manifests in deep learning models. Moreover, raindrops on the windshield, due to their refractive and scattering effects, can blur the entire image, creating considerable challenges for the model’s predictions. In the results from the left image, we can observe that YOLOP cannot accurately segment the drivable area due to reflections on the road surface. A similar situation is evident in the middle scene. YOLOP even struggles to distinguish between the car hood and the road. However, our model still delivers outstanding performance in this scenario. Especially for A-YOLOM(s), it is less affected by the reflections on the road surface. Our model also demonstrates superior lane line detection compared to YOLOP, as evidenced by the right image. However, in this scene, YOLOP outperforms us in detection tasks. As observed in the left image, YOLOP can detect more vehicles at a farther distance.
@@
In various weather conditions, our model consistently delivers more accurate and smoother segmentation results. Our model’s ability to detect distant and small targets falls slightly short compared to YOLOP in adverse weather conditions.
@@ IV-C Ablation Studies
Figure 6: Real Road Results
@@
Training method
Recall (%)
mAP50 (%)
mIoU (%)
Accuracy (%)
IoU (%)
YOLOM(n)
85.2
77.7
90.6
80.8
26.7
A-YOLOM(n)
85.3
78
90.5
81.3
28.2
YOLOM(s)
86.9
81.1
90.9
83.9
28.2
A-YOLOM(s)
86.9
81.1
91
84.9
28.8
In this subsection, we present two ablation studies to elucidate the effects of our proposed adaptive concatenation, as well as the impact of the segmentation neck and head on the overall model.
@@ IV-C1 Adaptive concatenation module
To assess the impact of our proposed adaptive concatenation module, we compare performance with and without this module. Specifically, YOLOM(n) and YOLOM(s) act as baselines, representing two distinct experiment groups with different backbones. We’ve carefully designed the segmentation neck structure for them to address the demands of particular segmentation tasks. However, both A-YOLOM(n) and A-YOLOM(s) include the adaptive concatenation module, so their segmentation neck structure is entirely identical. The results are presented in Table V. When comparing A-YOLOM(n) with YOLOM(n), we found comparable performance in detection and drivable area segmentation. However, both lane line accuracy and IoU showed significant improvements. Similarly, when comparing A-YOLOM(s) to YOLOM(s), we observe the same trend of improvement. These results indicate that our proposed adaptive concatenation module can proactively adapt to various tasks and are either competitive with or outperforms models that have been meticulously crafted based on experience.
@@ IV-C2 Multi-task model and segmentation structure
To evaluate the influence of the multi-task approach for each individual task, as well as our proposed segmentation neck and head structures, we compare performance and head parameters in YOLOv8(segda), YOLOv8(segll), YOLOv8(multi), and YOLOM(n) within the domain of segmentation tasks. YOLOM(n) have a similar backbone network with YOLOv8, making this comparison fair and make sense. The results are presented in Table VI. YOLOv8(segda) and YOLOv8(segll) implement the drivable area and lane line segmentation tasks separately. YOLOv8 (multi) is an integrated multi-task learning model that combines YOLOv8(segda), YOLOv8(segll), and YOLOv8(det) neck and head structure into one shared backbone. We observed there’s a significant improvement in performance. This indicated that multi-task learning can reciprocally enhance the performance of individual tasks.
@@
Model
Parameters
mIoU (%)
Accuracy (%)
IoU (%)
YOLOv8(segda)
1004275
78.1
YOLOv8(segll)
1004275
80.5
22.9
YOLOv8(multi)
2008550
84.2
81.7
24.3
YOLOM(n)
15880
90.6
80.8
26.7
Compared to YOLOv8(multi), YOLOM(n) has our carefully designed neck structure tailored for segmentation tasks. Additionally, it boasts a significantly more lightweight head structure, having only 0.008 times the complexity. This marked improvement stems from our unique head design, which solely relies on a sequence of convolutional layers, directly outputting a mask without the need for additional protos information. YOLOM(n) achieves impressive improved results in both mIoU and IoU for the drivable area and lane line tasks, respectively. Furthermore, we deliver competitive accuracy in the lane line task. These results evident our proposed neck and head innovations have achieved commendable results with minimal parameter and performance overhead.
@@ IV-D Real roads experiments
This section primarily discusses the experiments conducted in real road datasets. Specifically, we capture several videos using a dash camera and then convert them frame by frame into images to make predictions using our model and YOLOP. Each converted image dataset consists of 1428 images with a resolution of 1280x720. These images encompass three scenarios: highway, nighttime, and daytime. Figure 6 displays the results from the real road dataset. However, our model consistently maintains relatively stable performance, particularly for the A-YOLOM(s). In all tasks, A-YOLOM(s) outperforms YOLOP. Meanwhile, A-YOLOM(n) is slightly inferior only in detection tasks compared to YOLOP. This is crucial because ADS must be capable of operating in unfamiliar scenarios. The results on real roads demonstrate that our model can meet the actual needs of autonomous vehicles on real roads.
@@ V CONCLUSION
In this study, we primarily introduced an end-to-end lightweight multi-task model designed for real-time autonomous driving applications. The strength of a multi-task model lies in its capacity for tasks to implicitly benefit from and enhance each other, culminating in holistic performance improvements. To further bolster the model’s capabilities, we integrated an adaptive concatenation module, enhancing its flexibility. When compared to other state-of-the-art real-time multi-task methods using the BDD100k dataset, our model not only demonstrated superior visualization results but also showcased a faster processing speed. This speed gain is attributed to our streamlined segmentation head design, which optimizes both parameters and inference time. Moreover, real-road dataset evaluations underline our model’s robustness in novel environments, positioning it ahead of the competition. In the future, we aspire to encapsulate more autonomous driving tasks within our model and optimize it for edge device deployments.

### v4

@@ IV-B2 Multi-task Result
This part presents the results of multi-task experiment, encompassing object detection, drivable area segmentation, and lane line segmentation.
We follow the YOLOP settings in the detection task, combining the car, bus, truck, and train into a “vehicle” classification. The comparison results are shown in Table II. Based on the quantitative results, both our models achieve the best performance in terms of mAP50. This indicates that our model’s accuracy is excellent in predicting detected targets. Especially when compared to MultiNet and Faster R-CNN. Additionally, A-YOLOM(n) has a complexity level for its backbone, detection neck, and head that is comparable to YOLOv8n(det). However, A-YOLOM(n) exhibits significant improvement over YOLOv8n(det) in both recall and mAP50. This demonstrates that in multi-task learning, various tasks can implicitly assist and further enhance the performance of individual tasks. On another note, YOLOv8n(det) outperforms YOLOv5s because the ’s’ scale model usually has a more complex backbone to enhance performance. One issue with our model is the unsatisfactory recall. YOLOP and DLT-Net achieve better recall than our model. We believe this is due to the high weight of bounding box loss in $\mathcal{L}_{det}$. This means our model is more conservative and focuses on achieving a higher mAP at the expense of recall performance. We believe that mAP50 better reflects the comprehensive performance in the detection task. Besides, our model achieves real-time performance, while YOLOP cannot when the batch size is set to 1 according to Table I.
@@
Model Recall (%) mAP50 (%)
MultiNet 81.3 60.2
DLT-Net 89.4 68.4
Faster R-CNN 81.2 64.9
YOLOv5s 86.8 77.2
YOLOv8n(det) 82.2 75.1
YOLOP 88.6 76.5
A-YOLOM(n) 85.3 78.0
A-YOLOM(s) 86.9 81.1
For the drivable area segmentation task, Table III provides the quantitative results. Our model achieves the second and third-best performances in terms of mIoU. YOLOP outperforms our model because they customized the loss function for the segmentation task. In terms of mIoU, A-YOLOM(n) and A-YOLOM(s) trail YOLOP by 1% and 0.5% respectively. We believe this result is acceptable. While our model might exhibit a slight sacrifice in performance, it is more flexible and faster. Additionally, our model performs much better than other models, such as YOLOv8n(seg), MultiNet, and PSPNet. It’s worth noting that while YOLOv8n(seg) is faster than our model, its deployment cost is higher, and its performance is significantly inferior to ours.
@@
Model mIoU (%)
MultiNet 71.6
DLT-Net 72.1
PSPNet 89.6
YOLOv8n(seg) 78.1
YOLOP 91.6
A-YOLOM(n) 90.5
A-YOLOM(s) 91.0
For the lane line segmentation task, Table IV provides the quantitative results. A-YOLOM(s) achieves the best performance in terms of both accuracy and IoU. Specifically, A-YOLOM(s) delivers competitive results in accuracy and higher 2.3% in IoU compared to YOLOP. It’s worth noting that our model maintains the same structure and loss function across all segmentation tasks, eliminating the need for adjustments when encountering a new segmentation task. Additionally, YOLOv8 (seg) is slightly inferior to ours. Based on results in Table III and Table IV, we believe that our proposed neck and head are better suited for the segmentation task compared to YOLOv8 (seg). Since we use balance accuracy for evaluation, Enet, SCNN, and ENet-SAD employ different accuracy calculation approaches. Therefore, we cannot directly compare our accuracy with theirs. However, our results in IoU metrics are significantly better than theirs. This alone sufficiently demonstrates the superior performance of our model compared to theirs.
@@
Model Accuracy (%) IoU (%)
Enet N/A 14.64
SCNN N/A 15.84
ENet-SAD N/A 16.02
YOLOv8n(seg) 80.5 22.9
YOLOP 84.8 26.5
A-YOLOM(n) 81.3 28.2
A-YOLOM(s) 84.9 28.8
@@ IV-B3 Visualization
This section presents a visual comparison between YOLOP and our model. We evaluate performance not only in favourable weather conditions but also under adverse conditions, including strong sunlight, nighttime, rain, and snow. We will analyze each scene one by one.
Figure 3 displays the results from a sunny day. As we know, strong sunlight can affect a driver’s vision. Similarly, it impacts image acquisition from the camera, further affecting the model’s performance. This is a challenge for the drivable area and lane line segmentation. In this challenging scenario, our model outshines YOLOP. Specifically, under the conditions of intense sunlight (as seen in the left and right images), our model delivers accurate lane line predictions and provides smoother indications of drivable areas. Additionally, our model exhibits greater accuracy in detecting smaller and farther vehicles. As observed in the middle and right images, our model successfully detects vehicles located far in the distance, both on the road in opposite driving directions and near the house. We believe that our model outperformed YOLOP in this scene according to the visualization results.
@@
Figure 4 displays the results from night scenes. In nighttime scenes, the image quality decreases due to limited lighting and glare from oncoming vehicles. This is a challenge for the model to make accurate predictions. Under this challenge, our model consistently produces more accurate and smoother predictions for both lane lines and drivable areas. A-YOLOM(s) excels not only in the segmentation task but also in the detection task. Specifically, it surpasses A-YOLOM(n) by accurately detecting vehicles driving in the opposite direction, even amidst glare, and holds its own against YOLOP. Based on the detection results from the right image, YOLOP slightly outperforms our model in detecting distant vehicles at night. Nonetheless, our model produces significantly better segmentation results compared to YOLOP. Especially in the middle image, YOLOP mistakenly predicts the opposite lane as a drivable area. Such mispredictions are extremely hazardous for autonomous driving tasks.
@@
Figure 5 displays the results from a rainy day. Rain will increase road surface reflection, thereby impacting the driver’s judgment. This issue also manifests in deep learning models. Moreover, raindrops on the windshield, due to their refractive and scattering effects, can blur the entire image, creating considerable challenges for the model’s predictions. In the results from the left image, we can observe that YOLOP cannot accurately segment the drivable area due to reflections on the road surface. A similar situation is evident in the middle image. YOLOP even struggles to distinguish between the car hood and the road. However, our model still delivers outstanding performance in this scenario. Especially for A-YOLOM(s), it is less affected by the reflections on the road surface. Our model also demonstrates superior lane line detection compared to YOLOP, as evidenced by the right image. However, in this scene, YOLOP outperforms us in detection tasks. As observed in the left image, YOLOP can detect more vehicles at a farther distance.
@@
In various weather conditions, our model consistently delivers more accurate and smoother segmentation results. However, our model’s ability to detect distant and small targets falls slightly compared to YOLOP in adverse weather conditions.
@@ IV-C Ablation Studies
Figure 7: Real Road Results. Left image: nighttime. Middle image: daytime. Right image: highway.
@@
Training method Recall (%) mAP50 (%) mIoU (%) Accuracy (%) IoU (%)
YOLOM(n) 85.2 77.7 90.6 80.8 26.7
A-YOLOM(n) 85.3 78 90.5 81.3 28.2
YOLOM(s) 86.9 81.1 90.9 83.9 28.2
A-YOLOM(s) 86.9 81.1 91 84.9 28.8
In this subsection, we present an ablation study to validate the effectiveness of our adaptive concatenate module. Additionally, we evaluate the impact of the segmentation neck and head on the overall model performance.
@@ IV-C1 Adaptive concatenation module
To assess the impact of our adaptive concatenation module, we compare performance with and without this module. Specifically, YOLOM(n) and YOLOM(s) as baselines, representing two distinct experiment groups with different backbones. We have carefully designed the segmentation neck structure for them to address the demands of particular segmentation tasks. However, both A-YOLOM(n) and A-YOLOM(s) include the adaptive concatenation module. Their segmentation neck structure is entirely identical and without well-design. The results are presented in Table V. When comparing A-YOLOM(n) with YOLOM(n), we find comparable performance in detection and drivable area segmentation. However, both lane line accuracy and IoU show significant improvements. Similarly, when comparing A-YOLOM(s) to YOLOM(s), we observe the same trend of improvement. These results indicate that our adaptive concatenation module could adaptively concatenate features without manual design and achieve a similar or better performance than well-designed segmentation heads, further enhancing the model’s generality.
@@ IV-C2 Multi-task model and segmentation structure
To evaluate the influence of the multi-task approach for each individual task, as well as our proposed segmentation neck and head structures, we compare performance and head parameters in YOLOv8(segda), YOLOv8(segll), YOLOv8(multi), and YOLOM(n) within the domain of segmentation tasks. YOLOM(n) has a similar backbone network with YOLOv8, making this comparison fair and make sense. The results are presented in Table VI. YOLOv8(segda) and YOLOv8(segll) implement the drivable area and lane line segmentation tasks separately. YOLOv8(multi) is an integrated multi-task learning model that combines YOLOv8(segda), YOLOv8(segll), and YOLOv8(det) neck and head structure into one shared backbone. We observe there is a significant performance improvement. This indicates that multi-task learning can reciprocally enhance the performance of individual tasks.
@@
Model Parameters mIoU (%) Accuracy (%) IoU (%)
YOLOv8(segda) 1004275 78.1 - -
YOLOv8(segll) 1004275 - 80.5 22.9
YOLOv8(multi) 2008550 84.2 81.7 24.3
YOLOM(n) 15880 90.6 80.8 26.7
Compared to YOLOv8(multi), YOLOM(n) has our carefully designed neck structure tailored for segmentation tasks. Additionally, it boasts a significantly more lightweight head structure, having only 0.008 times the complexity. This significant improvement is due to our unique head design, which only relies on a series of convolutional layers, directly outputting a mask without the need for additional protos information. YOLOM(n) achieves impressive results in both mIoU and IoU for the drivable area and lane line tasks, respectively. Furthermore, we deliver competitive accuracy in the lane line task. These results evidence that our proposed neck and head innovations have achieved commendable results with minimal parameters and performance overhead.
@@ IV-D Real roads experiments
This section primarily discusses the experiments conducted in real road datasets. Specifically, we capture several videos using a dash camera and then convert them frame by frame into images to predict using our model and YOLOP. Each converted image dataset consists of 1428 images with a resolution of 1280x720. These images encompass three scenarios: highway, nighttime, and daytime. Figure 7 displays the results from the real road dataset. However, our model consistently maintains relatively stable performance, particularly for the A-YOLOM(s). In all tasks, A-YOLOM(s) outperforms YOLOP. Meanwhile, A-YOLOM(n) is slightly inferior only in detection tasks compared to YOLOP. ADS must be capable of operating smoothly in unfamiliar scenarios. This is crucial. The results on real roads demonstrate that our model can meet the needs of autonomous vehicles on real roads.
@@ V CONCLUSION
In this study, we primarily introduce an end-to-end lightweight multi-task model design for real-time autonomous driving applications. The advantages of the multi-task model lie in its capacity for each task to implicitly enhance the others, thereby further improving the performance of all tasks and enhancing the model’s deployability on edge devices. To improve the model’s generality, we integrate an adaptive concatenation module and propose unified loss functions for each type of task. This makes our model more flexible. When compared to other state-of-the-art real-time multi-task methods using the BDD100k dataset, our model not only demonstrates superior visualization results but also has a higher FPS. Moreover, real-road dataset evaluations prove our model’s robustness in novel environments, positioning it ahead of the competition. In the future, we aspire to encapsulate more autonomous driving tasks within our model and optimize it for edge device deployments.